# NestedMorph: Enhancing Deformable Medical Image Registration with Nested Attention Mechanisms

### v1

@@ Abstract
Deformable image registration is crucial for aligning medical images in a non-linear fashion across different modalities, allowing for precise spatial correspondence between varying anatomical structures. This paper presents NestedMorph, a novel network utilizing a Nested Attention Fusion approach to improve intra-subject deformable registration between T1-weighted (T1w) MRI and diffusion MRI (dMRI) data. NestedMorph integrates high-resolution spatial details from an encoder with semantic information from a decoder using a multi-scale framework, enhancing both local and global feature extraction. Our model notably outperforms existing methods, including CNN-based approaches like VoxelMorph, MIDIR, and CycleMorph, as well as Transformer-based models such as TransMorph and ViT-V-Net, and traditional techniques like NiftyReg and SyN. Evaluations on the HCP dataset demonstrate that NestedMorph achieves superior performance across key metrics, including SSIM, HD95, and SDlogJ, with the highest SSIM of 0.89, and the lowest HD95 of 2.5 and SDlogJ of 0.22. These results highlight NestedMorph’s ability to capture both local and global image features effectively, leading to superior registration performance. The promising outcomes of this study underscore NestedMorph’s potential to significantly advance deformable medical image registration, providing a robust framework for future research and clinical applications. The source code and our implementation are available at: https://bit.ly/3zdVqcg
@@ 1 Introduction
Deformable image registration is essential for medical imaging tasks as it aligns fixed and moving images by finding a non-linear mapping that minimizes their differences. This process establishes spatial correspondence through a continuous deformation field, which can be modeled in various ways, leading to numerous methods. Recent advancements in deformable image registration have been significantly enhanced by convolutional neural networks (CNNs), which have markedly improved both inference speed and accuracy compared to traditional methods. The field has seen a rapid rise in deep learning-based approaches. Recently, Vision Transformers (ViTs) have been integrated into deformable image registration to overcome the limitations of restricted receptive fields typical in CNN-based methods. However, Transformer-based techniques still face challenges in accurately modeling deformation fields. Here, we present a novel unsupervised hybrid CNN-Transformer based deep learning method for intra-subject registration, ensuring seamless alignment between T1w and dMRI templates for accurate anatomical and diffusion data correspondence.
dMRI of the brain induces distortions and partial volume effects from cerebrospinal fluid, necessitating a non-linear alignment approach. While signal averaging helps mitigate these effects, some studies opt for rigid body affine registration, which neglects nonlinear distortion [7]. To address this challenge, recent advancements in high-resolution dMRI have introduced nonlinear algorithms to improve alignment. Deep learning methods have also evolved, leveraging extensive datasets of paired or unpaired images to enhance deformable medical image registration, departing from traditional approaches [19].
Current research in medical image registration focuses on applying registration networks to diverse 3D images from public datasets but lacks comprehensive evaluation for intra-subject deformable registration between T1w and dMRI scans. While recent deep learning techniques like VoxelMorph [5] and CycleMorph [18] leverage CNNs for deformation fields, they struggle with long-range spatial relations [9]. In contrast, Transformer-based networks such as TransMorph [9] and ViT-V-Net [10] excel in understanding spatial correspondence, making them well-suited for intra-subject deformable registration between T1w and diffusion MRI scans [22]. Our proposed NestedMorph network is designed to improve intra-subject deformable registration between T1-weighted and dMRI modalities, and it includes a comparative analysis with state-of-the-art architectures as well as the proposed network itself. The key contributions of this study are summarized as follows
  • We present a Nested Attention Fusion network that integrates encoder high-resolution details with decoder semantic information, improving the alignment and accuracy of deformable registration.
@@
  • We compare our unsupervised method with leading CNN and Transformer-based approaches on the HCP dataset, showing superior performance in intra-subject deformable registration between dMRI and T1-weighted MRI data.
@@ 2 Related Works
Among the unsupervised methods, an unsupervised learning method for image registration that leverages a low-dimensional stochastic parameterization of deformation was proposed [21] . VoxelMorph [5] utilizes a CNN-based encoder-decoder architecture akin to UNet for computing deformation fields between moving and fixed image pairs, with a Spatial Transformer network warping the moving image iteratively. MIDIR [25] employs B-spline parameterization to generate diffeomorphic deformation fields, utilizing Mutual Information loss and a stationary velocity field (SVF) for smooth deformation. CycleMorph [18] incorporates two registration networks to compute displacement vector fields and introduces a cycle-consistent learning model for unsupervised registration while maintaining topology between images.
@@ 3 Proposed Method
Let $I_{f}$ and $I_{m}$ denote the fixed and moving images, which are the mean dMRI and T1w images, respectively. Let $\phi$ represent the deformation field that aligns the two images. The images $I_{f}$ and $I_{m}$ are concatenated and fed into the NestedMorph model. The goal here is to produce a deformation field $\phi$ that maps the grid of the moving image $I_{m}$ to the grid of the fixed image $I_{f}$, ensuring accurate alignment between the two. The encoder architecture employs a multi-level design that combines overlapping patch embeddings with dual attention mechanisms to extract and process features from 3D medical images. Initially, the input image is divided into overlapping patches and embedded into a lower-dimensional space through convolution operations. This embedding process reduces spatial resolution while preserving crucial information. The embedded patches are then processed through multiple stages, where efficient attention and channel attention mechanisms enhance the model’s ability to focus on both local and global features simultaneously. At each stage, the spatial dimensions are further reduced, and the feature richness is increased. The processed features are normalized and reshaped, resulting in a series of multi-scale feature maps. The decoder employs Dual Attention Enhanced Transformer (DAE-Former) blocks [3] to maintain long-range dependencies and structural integrity. The final layers incorporate a Large Kernel Attention (LKA) [4] module to handle local and global features, crucial for high-resolution image processing. The last layer outputs a 3-channel deformation field, used by the spatial transformer network (STN) to achieve precise image registration, ensuring alignment between $I_{f}$ and $I_{m}$.
@@ 3.1 Encoder
The NestedMorph model features an encoder block that integrates four encoder layers with dual Transformer mechanisms and a patch embedding module. Given the concatenated image of size $H\times W\times D\times 2$, the encoder processes the image through three stages. In the first stage, the image is divided into overlapping patches with a stride of 4, reducing the spatial dimensions to $\frac{H}{4}\times\frac{W}{4}\times\frac{D}{4}\times C_{1}$ channels. The output is then processed by the dual Transformer layers of the first encoder block, which apply Efficient Attention and Channel Attention mechanisms to capture both local and global dependencies.
@@ 3.1.1 Efficient and Channel Attention
where $\rho_{q}$ and $\rho_{k}$ are softmax normalization functions for queries and keys. Unlike dot-product attention, efficient attention bypasses the direct calculation of pairwise similarities. Instead, it employs keys represented as single-channel feature maps, which serve as weightings for all positions. These weightings are then used to aggregate value features through weighted summation, thereby constructing a global context vector. These global attention maps capture semantic aspects of the entire input feature, rather than just positional similarities. This approach reduces computational complexity while preserving high representational power.
@@ 3.1.2 Dual Attention Transformer Block
The dual attention transformer block integrates two main components: an Efficient Attention mechanism, followed by an Add & Norm step, and a Transpose Attention block that performs Channel Attention, succeeded by another Add & Norm step.
@@ 3.2 Decoder
In the decoder, the initial layers employ Dual Attention Enhanced Transformer (DAE-Former) blocks [3] to preserve long-range dependencies within lower-resolution images. These blocks are adept at integrating both spatial and channel attentions, ensuring that the image’s structural integrity is maintained while efficiently handling complex spatial interactions. The final layers of the decoder incorporate the Large Kernel Attention (LKA) [4] module, which excels in managing local and global features. This module addresses challenges associated with high-resolution image processing in transformers by enhancing the model’s ability to capture fine details and intricate patterns. Additionally, we employ a Nested Attention module that leverages skip connections to integrate features from various layers, enabling each decoder layer to access fine-grained spatial details crucial for generating accurate output masks. Specifically, we utilize cross-attention where the encoder’s output, denoted as $\mathbf{X_{2}}$, serves as the query input. This allows the decoder to focus on high-level features from the encoder. Meanwhile, the keys and values are derived from the output of the lower decoder layer, $\mathbf{X_{1}}$. This setup enables the decoder to successfully attend to and integrate both the detailed spatial information and the high-level contextual cues. The last layer outputs a deformation field $\phi$ with 3 channels, corresponding to the three spatial directions. This deformation field is then applied to a spatial transformer network for deformable registration, ensuring alignment between the moving $I_{m}$ and fixed $I_{f}$ images
@@ 3.2.1 Nested Attention Fusion Module
The Nested Attention Fusion Module, adapted from [20], processes decoder and encoder outputs ($\mathbf{X_{1}}$ and $\mathbf{X_{2}}$, respectively). The module utilizes multi-scale fusion to aggregate $\mathbf{X_{1}}$ and $\mathbf{X_{2}}$ in a semantically coherent manner, preparing the fused feature map for further processing. Spatial selection dynamically adjusts receptive fields to emphasize crucial features, while spatial interaction and cross-modulation enhance the feature maps with both local details and global context.
@@ 3.3 Spatial Transformer Network
Figure 3: Overall Framework of the Deformable Registration. This figure presents the complete NestedMorph registration pipeline, where the fixed and moving images are aligned using the NestedMorph network and a spatial attention network to generate a refined deformation field. The architecture integrates both local and global information, improving registration accuracy for medical image modalities such as T1-weighted MRI and dMRI.
@@ 4.2 Training Parameters
The NestedMorph model was implemented in Python using the PyTorch framework on an NVIDIA A100 GPU with 40 GB of memory. Training was conducted with a batch size of 4 using stochastic gradient descent, with a base learning rate of $1\times 10^{-4}$ and a weight decay of $3\times 10^{-5}$. The model was trained for 100 epochs using a composite loss function that computes the similarity between the deformed moving image and the fixed image, and imposes smoothness on the deformation field to regularize it [11].
@@ 4.4 Quantitative Results
From Table 1, it is evident that the proposed NestedMorph model outperforms all other models across the three evaluated metrics: SSIM, HD95, and SDlogJ. Specifically, NestedMorph achieves the highest SSIM of 0.89, indicating superior structural similarity and image quality after registration. It also demonstrates the lowest HD95 of 2.5 and SDlogJ of 0.22, reflecting better boundary alignment and reduced deformation irregularities, respectively. While models like TransMorph and CycleMorph show strong performance, with TransMorph coming close to NestedMorph in SSIM and having slightly higher HD95 and SDlogJ values, they do not surpass the NestedMorph model. Similarly, while NiftyReg and SyN improve upon the Affine method, they remain outperformed by the more advanced data-driven models. This comparison underscores the efficacy of integrating advanced modules into our proposed NestedMorph, enhancing its ability to deliver high-quality registrations with improved precision and reliability in medical images.
@@
Figure 4 presents a comparative analysis of several registration methods applied to medical images. The moving image ($I_{m}$) and the fixed image ($I_{f}$) are shown alongside the registered images generated by different methods. The top row displays the registered images, while the bottom row shows the corresponding binary difference images, which highlight the discrepancies between the registered and fixed images. The red boxes in the binary difference images highlight areas where the registration performance is easily noticeable. Each method’s performance is quantitatively assessed using SSIM, which is displayed above the binary difference images. The Affine method, although foundational, shows a moderate SSIM, reflecting its basic approach to alignment. NiftyReg and SyN show slight improvements in SSIM compared to the Affine method, but they still fall short of the performance achieved by deep learning-based approaches. In contrast, advanced techniques like MIDIR and VoxelMorph and CycleMorph improve the SSIM slightly, indicating better alignment and image fidelity. Overall, TransMorph and the proposed NestedMorph method achieve the highest SSIM scores, with NestedMorph slightly outperforming all others, signifying its superior performance in accurately aligning the moving image to the fixed image.
@@ 4.5.2 Attention Mechanism and Decoder Blocks
We also varied the number of DAE-Former and LKA blocks in the decoder. The optimal performance was observed with 2 DAE-Former and 2 LKA blocks, achieving a an SSIM of 0.840 and an HD95 of 2.4. This configuration balances feature relationship capture with contextual information. In contrast, other configurations with different combinations of DAE-Former and LKA blocks showed lower performance metrics. These results highlight the importance of balancing the number of blocks to achieve the best registration results. Overall, the combination of Efficient and Channel Attention mechanisms with 2 DAE-Former and 2 LKA blocks provides the most effective performance for the NestedMorph model.

### v3

@@ Abstract
Deformable image registration is crucial for aligning medical images in a nonlinear fashion across different modalities, allowing for precise spatial correspondence between varying anatomical structures. This paper presents NestedMorph, a novel network utilizing a Nested Attention Fusion approach to improve intra-subject deformable registration between T1-weighted (T1w) MRI and diffusion MRI (dMRI) data. NestedMorph integrates high-resolution spatial details from an encoder with semantic information from a decoder using a multi-scale framework, enhancing both local and global feature extraction. Our model notably outperforms existing methods, including CNN-based approaches like VoxelMorph, MIDIR, and CycleMorph, as well as Transformer-based models such as TransMorph and ViT-V-Net, and traditional techniques like NiftyReg and SyN. Evaluations using the HCP dataset demonstrate that NestedMorph achieves superior performance across key metrics, including SSIM, HD95, and SDlogJ, with the highest SSIM of 0.89, the lowest HD95 of 2.5 and SDlogJ of 0.22. These results highlight NestedMorph’s ability to capture both local and global image features effectively, leading to superior registration performance. The promising outcomes of this study underscore NestedMorph’s potential to significantly advance deformable medical image registration, providing a robust framework for future research and clinical applications. The source code and our implementation are available at: https://github.com/AS-Lab/Marthi-et-al-2024-NestedMorph-Deformable-Medical-Image-Registration
@@ 1 Introduction
Deformable image registration is essential for medical imaging tasks as it aligns fixed and moving images by finding a nonlinear mapping that minimizes their differences. This process establishes spatial correspondence through a continuous deformation field, which can be modeled in various ways, leading to numerous methods. Recent advancements in deformable image registration have been significantly enhanced by convolutional neural networks (CNNs), which have markedly improved both inference speed and accuracy compared to traditional methods. The field has seen a rapid rise in deep learning-based approaches. Recently, Vision Transformers (ViTs) have been integrated into deformable image registration to overcome the limitations of restricted receptive fields typical in CNN-based methods. However, Transformer-based techniques still face challenges in accurately modeling deformation fields. Here, we present a novel unsupervised hybrid CNN-Transformer based deep learning method for intra-subject registration, ensuring seamless alignment between T1w and dMRI templates for accurate anatomical and diffusion data correspondence.
dMRI of the brain induces distortions and partial volume effects from cerebrospinal fluid, necessitating a nonlinear alignment approach. While signal averaging helps mitigate these effects, some studies opt for rigid body affine registration, which neglects nonlinear distortion [7]. To address this challenge, recent advancements in high-resolution dMRI have introduced nonlinear algorithms to improve alignment. Deep learning methods have also evolved, leveraging extensive datasets of paired or unpaired images to enhance deformable medical image registration, departing from traditional approaches [19].
Current research in medical image registration focuses on applying registration networks to diverse 3D images from public datasets but lacks comprehensive evaluation for intra-subject deformable registration between T1w and dMRI scans. While recent deep learning techniques such as VoxelMorph [5] and CycleMorph [18] leverage CNNs for deformation fields, they struggle with long-range spatial relations [9]. In contrast, Transformer-based networks such as TransMorph [9] and ViT-V-Net [10] excel in understanding spatial correspondence, making them well-suited for intra-subject deformable registration between T1w and diffusion MRI scans [22]. Our proposed NestedMorph network is designed to improve intra-subject deformable registration between T1-weighted and dMRI modalities, and it includes a comparative analysis with state-of-the-art architectures as well as the proposed network itself. The key contributions of this study are summarized as follows
  • We present a Nested Attention Fusion network that integrates an encoder’s high-resolution details with a decoder’s semantic information, improving the alignment and accuracy of deformable registration.
@@
  • We compare our unsupervised method with leading CNN and Transformer-based approaches using the HCP dataset, showing superior performance in intra-subject deformable registration between dMRI and T1-weighted MRI data.
@@ 2 Related Works
Among the unsupervised methods, an unsupervised learning method for image registration that leverages a low-dimensional stochastic parameterization of deformation was proposed [21] . VoxelMorph [5] utilizes a CNN-based encoder-decoder architecture akin to UNet for computing deformation fields between moving and fixed image pairs, with a Spatial Transformer network warping the moving image iteratively. MIDIR [25] employs B-spline parameterization to generate diffeomorphic deformation fields, utilizing Mutual Information loss and a stationary velocity field (SVF) for smooth deformation. CycleMorph [18] incorporates two registration networks to compute the displacement vector fields and introduces a cycle-consistent learning model for unsupervised registration while maintaining topology between images.
@@ 3 Proposed Method
Let $I_{f}$ and $I_{m}$ denote the fixed and moving images, which are the mean dMRI and T1w images, respectively. Let $\phi$ represent the deformation field that aligns the two images. The images $I_{f}$ and $I_{m}$ are concatenated and fed into the NestedMorph model. The goal here is to produce a deformation field $\phi$ that maps the grid of the moving image $I_{m}$ to the grid of the fixed image $I_{f}$, ensuring accurate alignment between the two. The encoder architecture employs a multi-level design that combines overlapping patch embeddings with dual attention mechanisms to extract and process features from 3D medical images. Initially, the input image is divided into overlapping patches and embedded into a lower-dimensional space through convolution operations. This embedding process reduces spatial resolution while preserving crucial information. The embedded patches are then processed through multiple stages, where efficient attention and channel attention mechanisms enhance the model’s ability to focus on both local and global features simultaneously. At each stage, the spatial dimensions are further reduced, and the feature richness is increased. The processed features are normalized and reshaped, resulting in a series of multi-scale feature maps. The decoder employs a Dual Attention Enhanced Transformer (DAE-Former) blocks [3] to maintain long-range dependencies and structural integrity. The final layers incorporate a Large Kernel Attention (LKA) [4] module to handle local and global features, crucial for high-resolution image processing. The last layer outputs a 3-channel deformation field, used by the spatial transformer network (STN) to achieve precise image registration, ensuring alignment between $I_{f}$ and $I_{m}$. Figure 3 presents the overall framework of the deformable registration pipeline
@@ 3.1 Encoder
Figure 1 illustrates the NestedMorph architecture. The NestedMorph model features an encoder block that integrates four encoder layers with dual Transformer mechanisms and a patch embedding module. Given the concatenated image of size $H\times W\times D\times 2$, the encoder processes the image through three stages. In the first stage, the image is divided into overlapping patches with a stride of 4, reducing the spatial dimensions to $\frac{H}{4}\times\frac{W}{4}\times\frac{D}{4}\times C_{1}$ channels. The output is then processed by the dual Transformer layers of the first encoder block, which apply Efficient Attention and Channel Attention mechanisms to capture both local and global dependencies.
@@ 3.1.1 Efficient and Channel Attention
where $\rho_{q}$ and $\rho_{k}$ are softmax normalization functions for queries and keys. Unlike dot-product attention, efficient attention bypasses the direct calculation of pairwise similarities. Instead, it employs keys represented as single-channel feature maps, which serve as weightings for all positions. These weightings are then used to aggregate value features through weighted summation, thereby constructing a global context vector. These global attention maps capture semantic aspects of the entire input feature rather than just positional similarities. This approach reduces computational complexity while preserving high representational power.
@@ 3.1.2 Dual Attention Transformer Block
The dual attention transformer block integrates two main components: an Efficient Attention mechanism followed by an Add & Norm step, and a Transpose Attention block that performs Channel Attention, succeeded by another Add & Norm step.
@@ 3.2 Decoder
In the decoder, the initial layers employ Dual Attention Enhanced Transformer (DAE-Former) blocks [3] to preserve long-range dependencies within lower-resolution images. These blocks are adept at integrating both spatial and channel attention, ensuring that the image’s structural integrity is maintained while efficiently handling complex spatial interactions. The final layers of the decoder incorporate the Large Kernel Attention (LKA) [4] module, which excels in managing local and global features. This module addresses challenges associated with high-resolution image processing in transformers by enhancing the model’s ability to capture fine details and intricate patterns. Additionally, we employ a Nested Attention module that leverages skip connections to integrate features from various layers, enabling each decoder layer to access fine-grained spatial details crucial for generating accurate output masks. Specifically, we utilize cross-attention where the encoder’s output, denoted as $\mathbf{X_{2}}$, serves as the query input. This allows the decoder to focus on high-level features from the encoder. Meanwhile, the keys and values are derived from the output of the lower decoder layer, $\mathbf{X_{1}}$. This setup enables the decoder to successfully attend to and integrate both the detailed spatial information and the high-level contextual cues. The last layer outputs a deformation field $\phi$ with 3 channels, corresponding to the three spatial directions. This deformation field is then applied to a spatial transformer network for deformable registration, ensuring alignment between the moving $I_{m}$ and fixed $I_{f}$ images
@@ 3.2.1 Nested Attention Fusion Module
The Nested Attention Fusion Module, adapted from [20], processes decoder and encoder outputs ($\mathbf{X_{1}}$ and $\mathbf{X_{2}}$, respectively). The module utilizes multi-scale fusion to aggregate $\mathbf{X_{1}}$ and $\mathbf{X_{2}}$ in a semantically coherent manner, preparing the fused feature map for further processing. Spatial selection dynamically adjusts receptive fields to emphasize crucial features, while spatial interaction and cross modulation enhance the feature maps with both local details and global context. Figure 2 shows the Nested Attention Fusion Module.
@@ 3.3 Spatial Transformer Network
Figure 3: Overall Framework of the Deformable Registration. This figure presents the complete NestedMorph registration pipeline, where the fixed and moving images are aligned using the NestedMorph network and a spatial transformer network to generate a refined deformation field. The architecture integrates both local and global information, improving registration accuracy for medical image modalities such as T1-weighted MRI and dMRI.
@@ 4.2 Training Parameters
The NestedMorph model was implemented in Python using the PyTorch framework on an NVIDIA A100 GPU with 40 GB of memory. The training was conducted with a batch size of 4 using stochastic gradient descent, with a base learning rate of $1\times 10^{-4}$ and a weight decay of $3\times 10^{-5}$. The model was trained for 100 epochs using a composite loss function that computes the similarity between the deformed moving image and the fixed image, and imposes smoothness on the deformation field to regularize it [11].
@@ 4.4 Quantitative Results
From Table 1, it is evident that the proposed NestedMorph model outperforms all other models across the three evaluated metrics: SSIM, HD95, and SDlogJ. Specifically, NestedMorph achieves the highest SSIM of 0.89, indicating superior structural similarity and image quality after registration. It also demonstrates the lowest HD95 of 2.5 and SDlogJ of 0.22, reflecting better boundary alignment and reduced deformation irregularities, respectively. While models like TransMorph and CycleMorph show strong performance, with TransMorph coming close to NestedMorph in SSIM and having slightly higher HD95 and SDlogJ values, they do not surpass the NestedMorph model. Similarly, while NiftyReg and SyN improve upon the Affine method, they remain outperformed by the more advanced data-driven models. This comparison underscores the efficacy of integrating advanced modules into our proposed NestedMorph, thus enhancing its ability to deliver high-quality registrations with improved precision and reliability in medical images.
@@
Figure 4 presents a comparative analysis of several registration methods applied to medical images. The moving image ($I_{m}$) and the fixed image ($I_{f}$) are shown alongside the registered images generated by different methods. The top row displays the registered images, while the bottom row shows the corresponding binary difference images, which highlight the discrepancies between the registered and fixed images. The red boxes in the binary difference images highlight areas where the registration performance is easily noticeable. Each method’s performance is quantitatively assessed using SSIM, which is displayed above the binary difference images. The Affine method, although foundational, shows a moderate SSIM, reflecting its basic approach to alignment. NiftyReg and SyN show slight improvements in SSIM compared to the Affine method, but they still fall short of the performance achieved by deep learning-based approaches. In contrast, advanced techniques such as MIDIR and VoxelMorph and CycleMorph improve the SSIM slightly, indicating better alignment and image fidelity. Overall, TransMorph and the proposed NestedMorph method achieve the highest SSIM scores, with NestedMorph slightly outperforming all others, signifying its superior performance in accurately aligning the moving image to the fixed image.
@@ 4.5.2 Attention Mechanism and Decoder Blocks
We also varied the number of DAE-Former and LKA blocks in the decoder. The optimal performance was observed with 2 DAE-Former and 2 LKA blocks, achieving an SSIM of 0.840 and an HD95 of 2.4. This configuration balances feature relationship capture with contextual information. In contrast, other configurations with different combinations of DAE-Former and LKA blocks showed lower performance metrics. These results highlight the importance of balancing the number of blocks to achieve the best registration results. Overall, the combination of Efficient and Channel Attention mechanisms with 2 DAE-Former and 2 LKA blocks provides the most effective performance for the NestedMorph model.